7. Determine the returns to scale of the following productions functions


A. Q = 0.5L0.5K0.3


B. Q = 10L2 + 5K2


C. Q = 5L0.4K0.6


8. Based on the following table answer the questions that follow


labor Total product Marginal product Average product


0 0


1 3


2 8


3 12


4 14


5 14


6 12



A. Compute marginal product and average product of labor


B. Show the three stages of production based on the above data



1
Expert's answer
2022-02-01T14:01:03-0500

Production function: Y = F(K,L)

We augment the T to the factor t

"f(tL, tK) = t^{\u03b1 + \u03b2} + f(L,K)"

if α + β = 1, the f(L,K) exhibit Constant Returns to Scale

if α + β > 1, the f(L,K) exhibit Increasing Returns to Scale

if α + β < 1, the f(L,K) exhibit Decreasing Returns to Scale

i) Q = 2K + 3L + KL

Let K = 3 and L = 3

Q = 6 + 9 + 9 = 24

Let’s double the input, L = 6 and K = 6

Q = 12 + 18 +36 = 66

We observe that output got more than double by doubling the input values.

So, Q exhibits Increasing Returns to Scale.

ii) Q = 20K0.6L0.5

Augmenting Q by factor λ

Q(λK, λL) = 30 (λL)0.6(λK)0.5 = λ^1.1 \times Q(K,L)

Since, α + β = 1.1 > 1

So, Q exhibits Increasing Returns to Scale.

iii) Q = 100 + 3K + 2L

Let K = 2 and L = 2

Q = 100 + 6 + 4 = 110

Let's double the inputs, L = 4 and K = 4

Q = 100 + 12 + 8 = 120

We observe that output got less than double by doubling the input values.

So, Q exhibits Decreasing Returns to Scale.

iv) Q = 5KaLb, Where a+ b = 1

Augmenting Q by factor λ

Q(λK, λL) = 5(λL)a(λK)b = "λ^1\times Q(K,L)"

Since, α + β = 1

So, Q exhibits Constant Returns to Scale.

v) Q = K/L

Augmenting Q by factor λ

Q(λK, λL) = (λK / λL) = (K/L)

So, the production function exhibits Constant Returns to Scale